| Operation Summary for February 6, 2012 - February 8, 2012 |
|
The day shift began with Operations monitoring the Main Injector (MI) Lambertson magnet bake. During the day shift, Operators prepared Booster for an access at 8:17 AM to replace the power amplifier for a RF station (BRF18). Operators reported at 8:31 AM that they had searched and secured the ILC Test Facility (NML). At 8:32 AM, a Linac RF station (LRF1) tripped off and would not reset; Linac technicians investigated. Power Supply technicians reported at 9:25 AM that they had completed their work on the Switchyard (SW) power supplies at sectors F3 and F4. Booster came out of access at 10:52 AM. MI personnel reported that all tunnel work was completed; Operations continued to monitor the Lambertson bake. Operations established beam to MiniBooNE at 11:11 AM. Operators reported at 2:20 PM that MI had been ramping for MI-30 transformer commissioning. During the evening shift, Linac personnel reported at 5:39 AM that they would be running the I- Source at various energies from 700 to 750 KeV for studies. Operators had MI-65 searched and secured by 6:18 PM. Operators reported at 7:47 PM that they had turned on all the NuMI power supplies and were plotting them. At 11:49 PM, LRF3 tripped off and would not reset; Operators called in Linac personnel. Tuesday February 7 The midnight shift began with Linac off due to LRF3 problems; Operations had been sending beam to MiniBooNE. NuMI remained off due to the MI Lambertson bake. During the midnight shift, Linac resumed sending beam at 12:40 AM. Operators completed a Booster injection tune at 2:16 AM. Operators completed a 750 KeV line tune at 6:23 AM. At 6:31 AM, Operators began preparing MI for an access. At 7:22 AM, the Booster Gradient Magnet Power Supply tripped off; Power Supply personnel investigated and brought it back online. Operators prepared F sector for an access to let Water Group personnel search for a LCW leak. At 7:33 AM, Neutron Therapy Facility (NTF) personnel prepared to treat two patients later in the morning. Operators began preparing the Transfer Hall at 7:42 AM for an access to let Water Group personnel search for a LCW leak. During the day shift, at 8:42 AM, a panel breaker for a Linac quadrupole magnet power supply (L:QPSC02) tripped off; it reset. At 9:09 AM, L:QPS104 tripped off its panel breaker; it reset. Operators reported at 11:16 AM that the Lambertson bake was done; now the magnets need to cool down. Operators prepared MiniBooNE for an access at 1:36 PM. Linac personnel began some destructive beam energy studies at 1:56 PM. Three Linac quadrupole magnet power supplies (L:QPS105 and 108, and L:QPSC02) tripped off; Operators investigated and turned them back on. LRF4 and 5 tripped off for the first of two times in four minutes; Linac technicians investigated and turned then on. During the evening shift, Operations reported at 4:41 PM that they had received permission to send beam to the returning FTBF Meson experiment T-1017. Operators reported at 6:13 PM that the temperature of the Lambertson magnets had dropped and techs were taking equipment out of the tunnel. MI personnel asked to have Water Group personnel turn on the LCW for the Lambertsons in about three hours. Operators searched and secured the MTest enclosures for T-1017 at 7:41 PM. At 7:48 PM, LRF4 tripped off; a Linac tech investigated and turned it back on. Operations established beam to MiniBooNE at 7:58 PM. Operators prepared MI for access at 8:40 PM to check on the Lambertson magnets and valve in the LCW; the access turned off beam to MiniBooNE. L:QPS105 and 108 tripped off at 10:07 PM; Operators turned them back on. Operations established beam to MiniBooNE at 11:09 PM. By the end of shift, Operations and MI personnel had established beam to the MI dump. Wednesday February 8 The midnight shift began with a faulty MI ramp card and with MiniBooNE taking beam. During the midnight shift, Operators reported at 12:40 AM that they had replaced the MI ramp card. Operations reported that they had established beam to MI at 1:07 AM, but were intentionally aborting the beam every half-second while a MI person helped tune up injection. Operations established beam to NuMI around 1:30 AM. Operators reported at 1:45 AM that they had called in a Controls person to resolve problems with various autotune programs. At 2:05 AM, Operators reported that the vacuum at MI-40 would take some time to recover. Operators reported at 3:52 AM that they were having problems with MECAR aborts that didn’t seem to have a cause; a MI person had Operators turn on the automatic learning. Operations loaded protons into the Recycler at 4:46 AM. Operators prepared the Meson M01-M05 enclosures and MTest at 5 AM for Interlock personnel to conduct safety system tests. At 6:11 AM, LRF4 tripped off; it reset. A Controls person reported at the end of shift that the MI-8 autotune got confused because of new console name change; he fixed it. Operations Coordinator (OpsCo) Plans for Today The OpsCo agrees with plans by Booster personnel to conduct aperture scan studies and by Muon Rings personnel to conduct machine studies. MI personnel will continue to work on intensities and efficiencies. |
|
More Information Comments and Suggestions |
| last modified 02/08/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |